OpenStudy (anonymous):

solve the equation 43=7w-15-5w

OpenStudy (anonymous):

simplify the right side... what do you get?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

the right side is 2? so 43=2 ?

OpenStudy (anonymous):

43 = 2w - 15 58 = 2w w = 29

OpenStudy (anonymous):

43=2w-15 43+15=2w 58=2w 58/2=2w/2w 29=w w=29